# Self-Supervised Learning based on Transformed Image Reconstruction for Equivariance-Coherent Feature Representation

### v1

@@ 4.2 Synthetic tasks
Cross atten recon [36] 0.893 $\pm$ 0.004 0.921 $\pm$ 0.006 0.823 $\pm$ 0.030 0.875 $\pm$ 0.005 0.878
Ours(SIE, rot, rot) 0.9975 $\pm$ 0.0005 0.9073 $\pm$ 0.0021 0.9310 $\pm$ 0.0020 0.9810 $\pm$ 0.0010 0.954
Ours(SIE, all, rot) 0.9983 $\pm$ 0.0005 0.9231 $\pm$ 0.0005 0.9689 $\pm$ 0.0099 0.9801 $\pm$ 0.0004 0.968
Ours(SIE, all, color) 0.9891 $\pm$ 0.0019 0.9373 $\pm$ 0.0013 0.9700 $\pm$ 0.0067 0.9699 $\pm$ 0.0022 0.967
Ours(SIE, all, blur) 0.9981 $\pm$ 0.0001 0.9154 $\pm$ 0.0006 0.9392 $\pm$ 0.0106 0.9774 $\pm$ 0.0007 0.958
Ours(SIE, all, trans) 0.9975 $\pm$ 0.0005 0.9288 $\pm$ 0.0012 0.9747 $\pm$ 0.0017 0.9830 $\pm$ 0.0004 0.971
Ours(SIE, all, se(2)) 0.9980 $\pm$ 0.0001 0.9158 $\pm$ 0.0005 0.9520 $\pm$ 0.0050 0.9740 $\pm$ 0.0001 0.960
Table 2: Comparison of $R^{2}$ values across different configurations for synthetic tasks. Text in brackets in the configuration names indicate (1) for SIE the training transformation (cmp. Table 1) and (2) for Ours the first entry indicates the invariant loss, the second the augmentation applied to the input image, the third the transformation of the reconstruction asked for in $L_{\text{recon}}$. We apply either rotation as the sole augmentation or a combination of all augmentations (rot or all) and infer intermediate transformations such as angle, color, blur, and translation. Bold values indicate the best overall, while underlined values represent the best among baseline methods.
@@
Following [36], we design synthetic tasks to evaluate the equivariant representations learned during pretraining, see Figure 2. These regression-based tasks assess the transformation parameters between original and transformed views. Following the evaluation metrics used in SIE [18], we formulate transformation parameters prediction as a regression problem. To quantify alignment between predicted and true values, we use the coefficient of determination ($R^{2}$)
$R^{2}=1-\frac{RSS}{TSS}$ (2)
where $RSS$ is the sum of squared residuals, and $TSS$ is the total sum of squares. A higher $R^{2}$ value indicates more accurate transformation predictions.
4.2.1 Comparison with other SIE-based methods
@@
For this experiment, we use ViT-S/16 for all methods. Our method (Ours) uses the SIE-configuration and the SIE invariant loss for $L_{\text{SSL}}$. We compare with SIE [18] and a closely related cross-attention-based reconstruction method [36], see Table 2. For SIE, the performance with single augmentations and prior knowledge is the best among the baseline methods, but for color jitter estimation, where all variants perform well. However, it generalizes poorly to other transformation evaluations (best mean 0.610). While cross-attention reconstruction [36] leads to much more balanced results (mean 0.878), its performance does not surpass SIE with single augmentations, but for color. For our method, we tested different combinations of augmentation and intermediate transformation. All of them show a strong average performance increase (mean from 0.954 to 0.971 compared to 0.878 for the cross attention method [36]). The best ones, i.e. Ours(SIE, all, rot) and Ours(SIE, all, trans), even outperform all competitors on all tasks individually.
4.2.2 Enhancing the transformation prediction of SOTA augmentation-based SSL methods
@@
We test pretrained state-of-the-art Vit-L/16 models on the same rotation and color jitter prediction task (as outlined in 2), see Table 3. The baseline iBOT [45] pretrained model is taken from the official repository. For fair comparison, numbers shown for the DINOv2 baseline model are for a model we pretrained from scratch, as it performed better than with the weights from the official repository. For color prediction, iBOT and DINOv2 perform on par with the smaller ViT-S/16 models from Table 2; we repeated the entries for SIE(rot), SIE(all), and Ours(SIE, all, rot), for convenience. However, DINOv2 and iBOT perform significantly worse on the rotation estimation task. Training them from scratch using our approach improves their performance to best-in-class for color prediction (Ours(iBOT, all, se(2)). For rotation prediction, improvements for iBOT are remarkably high from 0.238 to 0.937. DINOv2 performance is also improved, but much less and to a lower performance (0.84). Notably, using se(2) as intermediate transform for iBOT and DINOv2 works slightly better than rotation, in contrast to SIE, see Table 2.
Configuration or Method Rotation Color
SIE(rot) 0.990 0.867
SIE(all) 0.331 0.899
IBOT 0.238 0.913
DINOv2 0.774 0.910
Ours(SIE, all, rot) 0.997 ($\uparrow$0.007) 0.923 ($\uparrow$0.056)
Ours(iBOT, all, rot) 0.925 ($\uparrow$0.687) 0.934 ($\uparrow$0.021)
Ours(iBOT, all, se(2)) 0.937 ($\uparrow$0.699) 0.943 ($\uparrow$0.030)
Ours(DINOv2, all, rot) 0.812 ($\uparrow$0.038) 0.920 ($\uparrow$0.010)
Ours(DINOv2, all, se(2)) 0.840 ($\uparrow$0.066) 0.933 ($\uparrow$0.023)
@@
Table 3: Performance comparison on rotation and color prediction tasks with improvements of our methods (absolute values). See Table 2 for description of configuration names. Rows with the same color represent the baseline results used for comparison with our methods.

### v2

@@ 1 Introduction
Cross Atten Recon (2024a) 0.893 $\pm$ 0.004 0.921 $\pm$ 0.006 0.823 $\pm$ 0.030 0.875 $\pm$ 0.005 0.878
Ours(VICReg, rot, rot) 0.9975 $\pm$ 0.0005 0.9073 $\pm$ 0.0021 0.9310 $\pm$ 0.0020 0.9810 $\pm$ 0.0010 0.954
Ours(VICReg, all, rot) 0.9983 $\pm$ 0.0005 0.9231 $\pm$ 0.0005 0.9689 $\pm$ 0.0099 0.9801 $\pm$ 0.0004 0.968
Ours(VICReg, all, color) 0.9891 $\pm$ 0.0019 0.9373 $\pm$ 0.0013 0.9700 $\pm$ 0.0067 0.9699 $\pm$ 0.0022 0.967
Ours(VICReg, all, blur) 0.9981 $\pm$ 0.0001 0.9154 $\pm$ 0.0006 0.9392 $\pm$ 0.0106 0.9774 $\pm$ 0.0007 0.958
Ours(VICReg, all, trans) 0.9975 $\pm$ 0.0005 0.9288 $\pm$ 0.0012 0.9747 $\pm$ 0.0017 0.9830 $\pm$ 0.0004 0.971
Ours(VICReg, all, SE(2)) 0.9980 $\pm$ 0.0001 0.9158 $\pm$ 0.0005 0.9520 $\pm$ 0.0050 0.9740 $\pm$ 0.0001 0.960
Table 1: Comparison of $R^{2}$ values across different configurations for synthetic tasks. Configuration naming: SIE methods show training transformation in brackets. Our methods use format: Ours(SSL loss, augmentation, reconstruction target) where ‘SSL loss’ is the $\mathcal{L}_{SSL}$ function used, ‘augmentation’ is the DINOv2 common augmentation pipeline (rot = rotation only, all = all common augmentations), and ‘reconstruction target’ is the target transformation for $\mathcal{L}_{\text{recon}}$. Bold values indicate overall best. We use VICReg (Bardes et al. 2022) as the invariance loss, consistent with SIE’s approach.
  • New task for joint invariant-equivariant learning: We introduce a reconstruction-based task that learns generalised equivariant features, relaxing restrictive assumptions about linear equivariant mappings.
  • Strong empirical validation on synthetic benchmarks: Our method achieves superior performance on all synthetic equivariance tasks from (Wang et al. 2024a), demonstrating clear advantages over existing approaches including SIE.
  • Consistent improvements across diverse real-world tasks: Compared to strong baselines (iBOT (Zhou et al. 2022), DINOv2 (Oquab et al. 2023)), our approach improves performance on most evaluation tasks while maintaining competitive results on invariance benchmarks.
  • General framework compatible with existing SSL methods: Our intermediate reconstruction approach can be integrated with various SSL frameworks, allowing to enhance existing methods with equivariant capabilities.
2 Related Work
2.1 Equivariant Neural Networks
Since the early days of neural networks research, the exploration of symmetries in the data has played a significant role, reduced model complexity, and improved inference quality of models (Fukushima 1980). One might argue that without convolutional neural networks, inherently implementing approximately translational equivariance, computer vision models could not have made this progress in the field. However, built-in permutation equivariance in transformer architectures has also been the object of study (Xu et al. 2023). Equivariance in deep learning can be split in two sub-categories: studies on models that inherit built-in equivariance (Cohen and Welling 2016; Jenner and Weiler 2022) or models that gain this property by experience (Xiao et al. 2020; Dangovski et al. 2021; Wang et al. 2024b).
2.2 Self-Supervised Learning
State-of-the-art SSL methods learn feature representations by automatically labelling non-labeled data and applying supervised learning techniques. The assumption is that the learned feature representation is comprehensive enough to be used later in other tasks, denoted as downstream tasks. A large variety of different SSL methods have been proposed (He et al. 2020; Chen et al. 2020a; He et al. 2022a; Chen et al. 2020b; Zbontar et al. 2021; Caron et al. 2021a; Bardes et al. 2022; Lehner et al. 2023; Xie et al. 2024; Oquab et al. 2023). For an overview, we refer to (Balestriero et al. 2023).
For our purpose it is relevant, how models react to different transformations in the input space, i.e. if they maintain the information in the feature representation or if this information is suppressed. Older SSL methods proposed auxiliary tasks such as Jigsaw puzzle (Noroozi and Favaro 2016) or rotation estimation (Gidaris et al. 2018) that foster equivariance properties as the transformation properties need to be maintained in feature space. These methods have been overtaken by matching-type methods. They present different versions of the same semantic content to the model and motivate it to map them to nearby points in feature space. Semantically different images are mapped to points far apart. This is achieved either by contrastive learning approaches or by regularisation techniques like de-correlation (Zbontar et al. 2021; Bardes et al. 2022) or teacher-student architectures (Zhou et al. 2022; Oquab et al. 2023). Consequently, these matching-type methods learn feature representations that suppress the differences between the versions.
Another state-of-the-art SSL branche includes mask-based approaches that remove parts of the input image and reconstruct them or a transformed version of them, either in the original image space (He et al. 2022a; Bandara et al. 2022), or in feature space (Assran et al. 2023). Contrastive learning has been combined with masked approaches, but only pixel-accurate translation has been applied as augmentation (Huang et al. 2022). As these methods do not, apart from masking or cropping, rely on other transformations, they are by construction more open for equivariance. Our approach is closely related to SIE (Garrido et al. 2023b), combining the matching approach with an explicit model of transformations applied in the input space. Our approach can be seen as an extension not requiring knowledge about the transformation parameters.
In contrast to SIE, our approach reaches state-of-the-art results.
2.3 Equivariance vs. Invariance
A function $f$ is denoted as equivariant with respect to a transformation $t$ in the input space and a corresponding transformation $\hat{t}$ in the output space, if the function commutes with the transformations, i.e. $\hat{t}(f(x))=f(t(x))$. Here, $f$ is a deep learning model, input space is the space of images or videos, and output space is the feature space.
The definition includes the identity transformation in the output space such that invariance is always also equivariance. However, in the computer vision literature e.g. (Xiao et al. 2020; Dangovski et al. 2021; Devillers and Lefort 2022; Garrido et al. 2023a; Park et al. 2022; Gupta et al. 2023; Wang et al. 2024b), as we do in this paper, invariance is often opposed to equivariance to stress that all information about the transformation in the input space is still retrievable from the output space. The term equivariance is usually considered for a set of transformations, i.e. the function is said to be equivariant with respect to this set of transformations. Moreover, the set of transformations is considered a group transformation or, even stronger, a group representation of the transformation, i.e. a linear map, in the input and output space. However, not all transformations applied in computer vision can be modelled by group transformations like elastic distortions, crop-resize operations, or non-affine perspective transformations. In addition, transformations that can theoretically be formulated as group transformations might lose the corresponding group properties in practice. E.g. an image rotated around a general angle cannot be rotated back as parts of the image get lost during the forward transformation. In this paper, we do not restrict the model to be equivariant with transformations that belong to a certain further structure. We argue that equivariance is no value in itself but should serve as a property to help to learn a feature representation that contains all the information necessary for all kinds of downstream tasks. We do not restrict the transformation to form a group or require them to be linear in the feature space. Instead we motivate the model to maintain the information that is necessary to maintain input output relation such that transformed images in the input space can be retrieved by the representation. It shall be irrelevant if the transformation forms a linear transformation, a general group transformation, or even if the definition of equivariance is only approximately fulfilled in the feature representation. We denote this as equivariant-coherence.
3 Method
Our approach extends SSL frameworks with an auxiliary reconstruction task that learns equivariance-coherent features. The key innovation is intermediate reconstruction: rather than learning from just the original and final transformed images, we supervise the model to reconstruct images at multiple points along transformation trajectories. This design naturally encourages the model to retain information necessary to perform the considered transformations. Given transformation $g_{\theta}$ defined by a parameter vector $\theta=[\varphi;t_{x};t_{y};$...] of continuous parameters, we define $K$ intermediate transformations $g_{\theta_{1}},g_{\theta_{2}},\ldots,g_{\theta_{K}}$ where $\theta_{k}:=\frac{k\theta}{K+1}$ for $k\in\{1,2,\ldots,K\}$ are parameter vectors consisting of equidistant parameters values between no transformation besides the base transformation and the additional transformation defined by $\theta$. In a first step to generate views as input for our joint embedding SSL approach a first view $v_{1}=\mathcal{A}_{1}(I)$ is generated by means of a set $\mathcal{A}_{1}$ of augmentation transformations. The second view $v_{2}$ is generated by a second set $\mathcal{A}_{2}$ of augmentation transformations $u:=\mathcal{A}_{2}(I)$ and, in contrast to the first view, it undergoes a sequence of additional transformations (the transformations are listed in Table 2)
$u\to g_{\theta_{1}}(u)\to g_{\theta_{2}}(u)\to\cdots\to g_{\theta_{K}}(u)\to g_{\theta}(u)$ (1)
where the end of the sequence constitutes the second view $v_{2}:=g_{\theta}(\mathcal{A}_{2}(I))$ in our joint embedding SSL method.
Intermediate images $u_{k}:=g_{\theta_{k}}(u)$ are to be reconstructed by the SSL method acting as anchor points shaping the geometry of the feature space. We empirically determine that $K=2$ yields optimal performance (Table 9).
3.1 Feature Splitting
In our proposed framework, depicted in Figure 1, we introduce transformation $g$ as an additional augmentation applied exclusively to $u\in\mathbb{R}^{B\times 3\times 224\times 224}$ where $B$ denotes the batch size. Both views, $v_{1}$ and $v_{2}$, are processed through encoder $f$, which operates either with shared weights or within a student-teacher framework depending on the SSL method used. In the student-teacher setup, the teacher network is updated using an exponential moving average (EMA) of the student’s parameters.
The resulting feature representations $z_{i}\in\mathbb{R}^{B\times N\times d_{\text{patch}}}$, $i\in\{1,2\}$, are split into two complementary components along the feature dimension, where $d_{\text{patch}}=d_{\text{inv}}+d_{\text{equi}}$, $N$ is the number of patches, and $d_{\text{patch}}$ is the patch embedding dimension:
  • Invariant features $z_{i}^{\text{inv}}\in\mathbb{R}^{B\times N\times d_{\text{inv}}}$ are supervised using standard SSL losses, e.g., the iBOT loss.
  • Equivariant features $z^{\text{equi}}_{i}\in\mathbb{R}^{B\times N\times d_{\text{equi}}}$ are used to reconstruct intermediate transformed images.
The dimension of the equivariant feature component, denoted as $d_{\text{equi}}$, can be adjusted as needed. We perform sensitivity analyses on $d_{\text{equi}}$ to examine its influence.
3.2 Intermediate Transformation Reconstruction
We employ $K$ independent decoders, $D_{1},D_{2},\ldots,D_{K}$, that operate on the concatenated equivariant features $z^{\text{equi}}_{1};z^{\text{equi}}_{2}]$ to reconstruct intermediate transformed versions of the input image. To reduce computational cost and emphasize the encoder’s role during pretraining, we deliberately design simple decoders consisting of a single linear layer followed by four convolutional layers (see Figure 1). Our objective is not to achieve high-fidelity reconstruction, but rather to provide sufficient supervisory signal for the encoder $f$ to learn effective equivariant representations. The lightweight decoder design ensures that the primary learning occurs in the encoder without interference from complex reconstruction architectures. Each decoder produces reconstructions:
$\hat{u}_{k}=D_{k}([z^{\text{equi}}_{1};z^{\text{equi}}_{2}])\quad\text{for }k=1,2,\ldots,K$ (2)
The reconstruction loss $\mathcal{L}_{\text{recon}}$ is computed as the mean $L_{2}$ loss between each decoder’s prediction $\hat{u}_{k}$ and the corresponding ground truth intermediate images $u_{k}$
$\mathcal{L}_{\text{recon}}=\frac{1}{K}\sum_{k=1}^{K}\|\hat{u}_{k}-u_{k}\|_{2}^{2}$ (3)
This reconstruction objective is combined with the standard SSL loss $\mathcal{L}_{\text{SSL}}$ using a weighting hyperparameter $\lambda$:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{SSL}}(z_{1}^{\text{inv}},z_{2}^{\text{inv}})+\lambda\mathcal{L}_{\text{recon}}$ (4)
A sensitivity analysis for $\lambda$ and $d_{\text{equi}}$ is given in Section 4.5.
3.3 Transformation Types and Parameters
Transformations $g$ Parameters Mag. range
Rotation angle $\varphi$ [45,45][-45,45]
Color jittering brightness, contrast, [0.4,0.4][-0.4,0.4]
saturation $S$, hue $H$ [0.1,0.1][-0.1,0.1]
Gaussian blur radius $\sigma$ $[0.1,2]$
Translation displacement $t_{x},t_{y}$ [10,10][-10,10]
SE(2) angle $\varphi$ [45,45][-45,45]
displacement $t_{x},t_{y}$ [10,10][-10,10]
Table 2: Considered transformations to generate the intermediate transformed images used as targets of the auxiliary reconstruction task, as well as the second view for the joint embedding.
Additionally, we incorporate non-geometric transformations, including color jittering and Gaussian blur. Transformation parameter ranges are shown in Table 2.
4 Experimental Results
As our method builds on SIE (Garrido et al. 2023b), SIE is our first natural baseline with synthetic evaluation tasks designed to benefit from equivariant features. To evaluate the generalization of our method against state-of-the-art approaches, we go beyond using the invariant loss function $L_{\text{SSL}}$ as in SIE and explore integrating other methods. Specifically, we incorporate co-learning with iBOT (Zhou et al. 2022) and DINOv2 (Oquab et al. 2023), both augmentation-based techniques. They are also tested on the synthetic benchmark. Finally, we evaluate on a rich set of more realistic downstream tasks. We aim to enhance the baselines’ performance on equivariance-related tasks while preserving strong results on invariance-related benchmarks, e.g. ImageNet linear probing.
4.1 Implementation Details
Architecture
We use Vision Transformers (ViTs) (Dosovitskiy et al. 2020) with different configurations, specifically ViT-S/16 and ViT-L/16, as backbones for experiments. We incorporate a linear head on top of the backbone as originally done by the baseline methods to accommodate different representation dimensions $d_{\text{patch}}$, i.e. 8192 for iBOT, 512 for SIE, and 2048 for DINOv2.
Afterwards, a portion $z^{\text{equi}}$ of these features $z$ is allocated for reconstruction. We do not introduce more features than the baseline methods.
Pretraining Setup
Our approach uses a baseline SSL loss $L_{\text{SSL}}$ in addition to our new component $L_{\text{recon}}$. Each of the three baseline methods come with distinct training setups. The common training configuration includes ImageNet-1K as the dataset, optimizer AdamW (Loshchilov and Hutter 2017), and a cosine-scheduled learning rate. For the SIE-based method, we apply their invariant loss as $L_{SSL}$, and pretrain ViT-S/16 for 800 epochs with a batch size of 2048. The base learning rate is set to $10^{-4}$ and is linearly scaled with the batch size $B$: $lr=10^{-4}\cdot B/256$. For the iBOT-based approach, we pretrain ViT-S/16 for 800 epochs and ViT-L/16 for 250 epochs, both with a batch size of 1024. The learning rate follows the same linear scaling strategy, with a base learning rate of 5e45e{-4}. For the DINOv2-based training, we train ViT-L/16 with 100 epochs with batch size of 2048. The base learning rate is 4e34e{-3} with warmup 10 epochs. In the pretraining stage, we weight the reconstruction loss introduced by the auxiliary task using the optimal coefficient $\lambda=1$, as determined by the sensitivity analysis in Section 4.5. For the equivariant feature dimension $d_{\textit{equiv}}$, we also select the hyperparameter based on this sensitivity analysis, adopting a default value of $2048$. For DINOv2, we retain the same proportional relationship between feature dimensions and therefore set $d_{\textit{equiv}}=512$. For VICReg, we follow SIE and evenly split the embedding dimension to determine $d_{\textit{equiv}}$.
Computation Cost
We conduct all pretraining experiments on the JUWELS Booster (Jülich Supercomputing Centre 2021). For VICReg, we adapt the method to a ViT-S/16 backbone and train on 4 nodes with a total of 16 A100 GPUs. Both iBOT and DINOv2 are pretrained on 16 nodes with 64 A100 GPUs. The computational cost of the different SSL methods, along with the additional cost introduced by our auxiliary tasks, is summarized in Table 3. Overall, our lightweight intermediate-transformation reconstruction task adds only a modest overhead to the baseline SSL methods.
Config. Backbone Runtime s/epoch Overhead
VICReg ViT-S/16 156 -
+ Ours ViT-S/16 179 14.7%
iBOT ViT-L/16 230 -
+ Ours ViT-L/16 249 8.3%
DINOv2 ViT-L/16 750 -
+ Ours ViT-L/16 804 7.2%
Table 3: Computational overhead, where Ours refer to Ours(all, SE(2)), with all base augmentations and SE(2) transformation.
Figure 2: Synthetic tasks for evaluating equivariant representations. Transformation ($g$) is applied to the original image II. Both the transformed and original images are processed through a pretrained encoder $f$. A lightweight MLP $h$ then predicts the parameters of the applied transformation.
4.2 Performance on Synthetic Tasks
@@
Following (Wang et al. 2024a), we design synthetic tasks to evaluate the equivariant representations learned during pretraining, see Figure 2. These regression-based tasks assess the transformation parameters between original and transformed views. Following the evaluation metrics used in SIE (Garrido et al. 2023b), we formulate transformation parameters prediction as a regression problem. To quantify alignment between predicted and true values, we use the coefficient of determination ($R^{2}=1-\frac{RSS}{TSS}$), where $RSS$ is the sum of squared residuals, and $TSS$ is the total sum of squares. A higher $R^{2}$ value indicates more accurate transformation predictions.
Comparison with Other SIE-Based Methods
@@
Here, we use ViT-S/16 for all tests. Our approach adopts the SIE configuration and leverages VICReg (Bardes et al. 2022) as $L_{\text{SSL}}$, consistent with SIE (Garrido et al. 2023b). We compare with SIE and a closely related cross-attention-based reconstruction method (Wang et al. 2024a), see Table 1. For SIE, the performance with single augmentations and prior knowledge is the best among the baseline methods, but for color jitter estimation, where all variants perform well. However, it generalizes poorly to other transformation evaluations (best mean $R^{2}=0.610$). Cross-attention reconstruction (Wang et al. 2024a) leads to much more balanced results (mean $R^{2}=0.878$). For our method, we tested different combinations of augmentation and intermediate transformation. All of them show a strong average performance increase (mean $R^{2}$ from 0.954 to 0.971).
The best ones, i.e. Ours(VICReg, all, rot) and Ours(VICReg, all, trans), outperform all competitors on all tasks individually, demonstrating the most versatile equivariant representations.
Enhancing Transformation Prediction of Augmentation-Based SSL Methods
@@
We test pretrained state-of-the-art ViT-L/16 models on the same rotation and color jitter prediction task (as outlined in 4.2), see Table 4. The baseline iBOT (Zhou et al. 2022) pretrained model is taken from the official repository. For fair comparison, numbers shown for the DINOv2 baseline model are for a model we pretrained from scratch, as it performed better than with the weights from the official repository. For color prediction, iBOT and DINOv2 perform on par with the smaller ViT-S/16 models from Table 1. However, DINOv2 and iBOT perform significantly worse on the rotation estimation task. Training them from scratch using our approach improves their performance to best-in-class for color prediction (Ours(iBOT, all, SE(2)). For rotation prediction, improvements for iBOT are remarkably high from 0.238 to 0.937. DINOv2 performance is improved less and to a lower performance (0.84). Notably, using SE(2) as intermediate transform for iBOT and DINOv2 works slightly better than rotation, w.r.t. SIE, see Table 1.
Configuration $R^{2}(\text{rot})$ $R^{2}(\text{color})$ Mean($R^{2}$)
SimCLR 0.288 0.575 0.432
+ Ours(all, rot) 0.637 0.712 0.675
+ Ours(all, SE(2)) 0.705 0.723 0.714
IBOT 0.238 0.913 0.576
+ Ours(all, rot) 0.925 0.934 0.930
+ Ours(all, SE(2)) 0.937 0.943 0.940
DINOv2 0.774 0.910 0.842
+ Ours(all, rot) 0.812 0.920 0.866
+ Ours(all, SE(2)) 0.840 0.933 0.887
@@
Table 4: Performance comparison on rotation and color prediction tasks with improvements of our methods (absolute values). See Table 1 for description of configuration names.